OpenStudy (anonymous):

Factor this expression completely. a^3y+1. Both the 3 and the y are exponents.

OpenStudy (anonymous):

\[a^{3y}+1=(a^y)^3+1^3=(a^y-1)(a^{2y}+a^{y}+1)\]